OpenStudy (anonymous):

4 + (-5) + 6 -1+(-8)+3 -9+4+(-9) 7+(-9)+(-3) -15+(-10)+(-4) -6+20+8 11+(-14)+3 -13+2+(-5) 16+16+(-7) -4+(-8)+(-3) -6+(-15)+4 -2+(-9)+(-4)+18 23-(-7)+(-3)+5 -11+6+(-11)+2 -1+4+8+(-17) 3+(-8)+13+4 -9+(-6)+30+(-2) 12+(-7)+5+(-15) 98+98+(-99)+(-99)

OpenStudy (unklerhaukus):

yes

OpenStudy (anonymous):

plz help!

OpenStudy (anonymous):

which type of calculation is that?

OpenStudy (anonymous):

integers

OpenStudy (anonymous):

each line is a separate problem

OpenStudy (unklerhaukus):

\[4 + (-5) + 6=4-5+6=5\]

OpenStudy (unklerhaukus):

\[1+(-8)+3=1-8+3=-4\]

OpenStudy (unklerhaukus):

Are you really having trouble with these?

OpenStudy (anonymous):

then your way of writing that one is not good.

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

you can do most of these problems with a number line for example 4+(−5)+6